# Cryoscope: A Cryogenic Infrared Survey Telescope

### v1

@@ 4.4 Gamma Ray Bursts
Observations of long-duration GRBs at low redshifts have established that the vast majority of them are associated with core-collapse stripped-envelope supernovae of high kinetic energy (around $10^{52}$ erg, e.g. one order of magnitude larger than classical supernovae, Mazzali et al., 2021). Moreover, the concomitant detection of gravitational waves from the binary neutron star coalescence event of 17 August 2017 and of a short-duration GRB (GRB170817A) has definitively proven the association of short-duration GRBs with binary neutron star mergers (Abbott et al., 2017b).
Recently, however, the dichotomy between long GRBs connected with supernovae and short GRBs associated with kilonovae has become blurred, due to detections of long GRBs without bright supernovae and associations with kilonovae instead (Rastinejad et al., 2022; Levan et al., 2024), and short GRBs associated with supernovae (Rossi et al., 2022).
A large field-of-view infrared camera with rapid reaction to transients could timely cover large error areas in the sky, such as those currently provided by the Fermi/GBM, Swift/BAT, SVOM/ECLAIR and WXT onboard Einstein Probe, and detect very early counterparts when their brightness and variability amplitude is at a maximum. Observations in K-dark are particularly relevant for very high-redshift events at $z>6$, and for those GRBs that occur deep within dusty regions (see Fig. 19). In both cases, the optical light is suppressed, and IR observations are needed to achieve detections. These observations will address critical aspects of GRB science within early observations, such as the interplay of initial emission components (internal versus external shocks; formation of forward and reverse shock; impact and interaction with circumburst medium). At late epochs, the late, non-relativistic transition can be followed, and the unobscured star-formation rate of the host galaxy estimated. Overall, the IR observations provided by Cryoscope will allow us to construct GRB afterglow samples with minimal biases due to extinction and redshift, factors that have strongly affected previous GRB observations.
Figure 19: Estimated magnitudes for GRB afterglows in K-dark magnitudes derived from the sample of Kann et al. (2011) assuming a spectral slope of $F_{\nu}$ $\sim 1/\nu$. These light curves are corrected for extinction, which gives us a more realistic view of what is to be expected in K-dark, where the effect of extinction is less significant. We plot the 5-sigma detection limits for 15 minute exposures with the Pathfinder and the full Cryoscope. Our goal is to aim for a start of the GRB observations within 30 minutes to an hour of the alert with Cryoscope Pathfinder, and within 5 minutes of the alert for the full Cryoscope.
@@ 4.5 Solar and Interstellar Comets and Asteroids
Small bodies (asteroids and comets) are the remnants of the blocks that accreted to form the planets 4.6 Gyrs ago. Their importance as witnesses of solar system history has emerged in past decades, but their current description is insufficient to guide theoretical works: current observations cannot discriminate between mutually excluding hypotheses (such as a densely populated or an empty primordial asteroid belt, for instance: Morbidelli et al., 2009; Raymond & Izidoro, 2017). Models may accurately reproduce the dynamical structure of the asteroid belt, but they diverge in their prediction of the distribution of composition. It is therefore crucial to identify which compositions were native to the asteroid belt between Mars and Jupiter and which were implanted later, and what were the implantation mechanisms.
Characterization of a large sample of asteroids can be achieved through broad-band spectro-photometry (DeMeo & Carry, 2013, 2014). In that respect, the upcoming Legacy Survey of Space and Time (LSST) of the Vera C. Rubin Observatory will play a major role (LSST Collaboration et al., 2009; Jones et al., 2009). However, with filters limited to the visible range only, LSST will not disentangle ambiguous spectra of widely different compositions, whose near-infrared colors differ (DeMeo et al., 2009; Carry, 2018; Mahlke et al., 2022). By acquiring photometry in K-band at multiple epochs/geometries over several years, Cryoscope Pathfinder will provide for the first time the absolute magnitude and phase function (Muinonen et al., 2010) of asteroids in the near-infrared, in a similar fashion as LSST in the visible (the only way to determine the colors of small bodies with LSST, see Mahlke et al., 2021).
The difference of absolute magnitude in the visible and near-infrared will break the spectral degeneracy inherent to only visible observations, providing the compositional distribution of asteroids. The obtention of phase functions in the near-infrared also opens up new prospects for compositional mapping (Mahlke et al., 2022): these functions are known to be wavelength-dependent (Sanchez et al., 2012) but are seldom acquired with a filter other than Johnson V.
@@
In addition to the characterization of a sample of asteroids, Cryoscope Pathfinder will be able to discover new asteroids and comets which happen to be in its field-of-view during any scientific observation. Owing to its location in Antarctica, Cryoscope’s footprint will be located at high ecliptic latitude, where a deficit of detection of small bodies has been reported (Mahlke et al., 2018; Carry et al., 2021). Generally, if an asteroid was observed during only a short, one-night arc and was not followed up afterwards, its orbit determination will be poor. Such objects might be lost completely, or until future recovery observations. These one-night arc observations will mostly populate the Isolated Tracklet File of the Minor Planet Center. Such observations might be linked with previous or future observations done by other facilities and will contribute to a pool of objects with well-established orbits. Objects which happen to be in the field-of-view for several consecutive nights might offer better prospects of orbit determination (typically, observations from three different nights are needed for robust orbit determination of a main-belt asteroid, see Holman et al., 2018; Le Montagner et al., 2023, and references therein). Such objects will also be better placed for detection by our synthetic tracking pipeline as more images can be stacked together, increasing the objects’ S/N and enabling probes of more distant populations of small solar system bodies fainter than 20 mag (Burdanov et al., 2023).
@@ 4.6 Stellar Microlensing
Gravitational microlensing can locate galactic populations of dark compact objects such as primordial black holes and wide-orbit or free-floating planets. In microlensing, a lens object intervenes across an observer’s line of sight to a distant source star. The light from the source is bent around the gravitational influence of the lens, and two source images are produced. These images are separated only by microarcseconds and cannot be resolved from the source PSF. However, as the lens moves along a trajectory, the images aggregate to temporarily amplify the source flux (Paczynski, 1986). A stellar microlensing event is recognizable as a symmetric peak in the flux measurements of a star on a timescale proportional to the lens mass. Less massive microlenses produce shorter microlensing events. A planetary-mass lens produces a microlensing event that spans hours, while a solar-mass lens can generate a signal duration of several weeks. The probability of detecting a microlensing event scales with the volume and density of a stellar field, the total observation time, and the detector’s technical sensitivity and overheads.
Exoplanets: The transit and radial velocity techniques excel at detecting massive exoplanets orbiting close to their host stars. Using these techniques to discover small rocky planets or those beyond the snowline poses a challenge; the resulting variations in the data are often near the signal-to-noise limit and cannot be detected. However, such ‘hidden’ exoplanets produce significant perturbations in binary microlensing events. Unbound, ‘free-floating’ planets (FFP) additionally produce microlensing events. It is expected that FFPs are the result of dynamical ejection mechanisms between stellar-planet and planet-planet systems (Mróz et al., 2019). FFP detections are essential for understanding the evolution of planetary systems and determining the ratio of ejected planets to bound ones (Bachelet et al., 2022). Modelling from the 9 years of MOA II data suggests that $19^{+23}_{-13}$ FFPs exist per star (Sumi et al., 2023). Microlensing provides a direct method for investigating the formation rates and distributions of distant, low-mass, or unbound exoplanets.
Primordial Black Holes: The Primordial Black Hole (PBH) is predicted to form from critically collapsing density perturbations in the pre-inflationary universe (Zel’dovich & Novikov, 1966; Hawking, 1971). Interest in PBH formation hypothesis and detection has rapidly increased over the last ten years, spurred by the discussions from Bird et al. (2016) regarding the possibility that PBHs may explain the unexpectedly large progenitor masses of the LIGO-Virgo mergers. Additionally, PBH detection efforts are often in tandem with dark matter detection, with a galactic halo population of PBHs forming a simple and tidy solution to dark matter in a $\Lambda$CDM universe.
[…]
Current & Upcoming Surveys: Current optical microlensing surveys (OGLE (Mróz et al., 2017), KMTNet (Jung et al., 2024; Yang et al., 2023), and MOA (Sum, 2011)) can strongly detect a population of icy exoplanets and FFPs, with hundreds of detections recorded. Near-infrared detectors have also successfully detected microlensing events towards regions of high dust extinction. The United Kingdom Infrared Telescope (UKIRT) Microlensing Survey proved the observational power of the H- and K-band in discerning obscured lensing signals with five events of durations spanning 7-30 days (Shvartzvald et al., 2017). Although not designed for microlensing detections, the VISTA Variables in the Via Lactea Survey (VVV) detected in excess of 600 microlenses along a 10 longitude strip in the Galactic center (Kondo et al., 2023). The sparse sampling rate of VVV favored events produced by massive lens objects. Kaczmarek et al. (2022) applied statistical modeling to the VVV candidates to evaluate two lenses as neutron stars. Soon, the PRime-focus Infrared Microlensing Experiment (PRIME) is expected to detect over 40 planets per year towards the inner Galactic Bulge in the H-band with a 1.45 deg2 field-of-view (Kondo et al., 2023).
The Nancy Grace Roman Space Telescope includes a dedicated microlensing survey, aiming to detect exoplanets at moderate to large distances (1-10 AU) from the host star, with over 1400 bound planets with mass greater than 0.1 $M_{\oplus}$, and more than 250 FFPs (Penny et al., 2019; Johnson et al., 2020). However, the composite seven Galactic Bulge fields at a 15-minute cadence narrow the survey’s focus to predominantly planetary and stellar class lenses (Bachelet et al., 2022). Additionally, no extragalactic fields have been formalized in the Roman microlensing program.
[…]
Most microlensing surveys employ a small field-of-view for predominantly Galactic Bulge science. A small field-of-view produces inconsistent sampling as multiple fields constantly cycle across the wider sky. Therefore, many microlensing events are incompletely sampled across the signal duration, favoring longer-duration events and more massive microlenses.
Cryoscope: Dust obscuration reduces the productive sightlines for searching for events in optical filters toward the Galactic disk and center (Shvartzvald et al., 2018). The disk has high stellar density and, consequently, large microlensing probability, but is riddled with high extinction levels. Imaging in the K-band enables Cryoscope to peer through the dust and image microlensing events in similar regions to the VVV and UKIRT surveys, albeit at a much faster cadence.
Cryoscope performs optimally in the key areas of improving microlensing detection. Cryoscope’s quarter-arcsecond seeing stability removes much of the atmospheric obstruction that commonly contaminates ground-based imaging of fast and low signal-to-noise microlensing events. The wide field-of-view of Cryoscrope can maximize the number of constantly monitored stars while reducing the overhead needed for cyclical slewing patterns. Using the LMC as an example, the 16.2 deg2 field-of-view of Cryoscope Pathfinder can cover the central bulge and peripheral fields within seven pointings. As a comparison example, DECam (3 deg2 field-of-view) would require 30 fields to cover the same LMC footprint. Therefore, for Cryoscope, fewer observation nights are required to perform a complete low-mass microlensing survey towards the LMC. Finally, multiple sightlines towards the Galactic Bulge, disk, and extragalactic fields (including LMC, SMC, and M31) are readily available for microlensing detection as a byproduct of a high-volume survey. Tracing regions of high stellar density will increase the probability of detecting multiple microlensing signals, providing better event statistics to distinguish between the kinematics of baryonic (planetary) and non-baryonic (PBH) microlenses.
@@ 4.7 Core Collapse Supernovae
Massive stars with $\gtrsim 8~{}M_{\odot}$ form an iron core towards the end of their life and die when this iron core exceeds the Chandrasekhar mass and collapses, sending out a shock wave that disrupts the star. The resulting core-collapse supernova (CCSN) forms a diverse group of transients whose properties are determined by the properties of the progenitor star. For instance, the amount of hydrogen left in the envelope determines the spectral type of the CCSN (Filippenko, 1997); the star’s mass at the time of explosion determines how long the light curve takes to rise (Khatami & Kasen, 2019); and the core mass of the star determines the strengths of different emission lines later in the nebular phase (e.g., Jerkstrand et al., 2014). While most of what we know about CCSNe as a class have been derived from optical observations, increasing IR observations in the past decade revealed several aspects of massive stellar explosions that had remained elusive. Cryoscope observations will help us probe the chemical evolution of CCSNe ejecta, reveal late-time power sources of CCSNe, and uncover a population of CCSNe hidden behind dust.
@@
While JWST observations are extremely powerful in tracking molecule and dust formation in CCSNe, the number of observations is still limited to a few nearby and notable objects (e.g., Hosseinzadeh et al., 2023; Shahbandeh et al., 2023). Cryoscope is unique as the only planned all-sky survey that covers wavelengths beyond 2 $\mu$m (unlike the DREAMS and WINTER surveys powered by InGaAs detectors). The K-dark band, unlike the more typical Ks, covers the bandhead of CO’s first overtone band at around 2.3 $\mu$m. K-dark observations will also be able to track the rising thermal continuum from newly formed dust grains in the ejecta of CCSNe. Unlike in previous targeted IR observations, Cryoscope will observe a statistically significant sample of CCSNe and measure the evolution of CO and dust formation to test chemical evolution models of CCSNe.

### v2

@@ 4.4 Gamma Ray Bursts
Cryoscope’s large field-of-view and high cadence in the $K_{dark}$ bandpass will be well-suited for detecting early counterparts to dust-enshrouded and high-redshift gamma ray bursts (GRBs).
@@
Observations of long-duration GRBs at low redshifts have established that the vast majority of them are associated with core-collapse stripped-envelope supernovae of high kinetic energy (around $10^{52}$ erg, e.g. one order of magnitude larger than classical supernovae, Mazzali et al., 2021). Moreover, the concomitant detection of gravitational waves from the binary neutron star coalescence event of 17 August 2017 and of a short-duration GRB (GRB170817A) has definitively proven the association of short-duration GRBs with binary neutron star mergers (Abbott et al., 2017b). Recently, however, the dichotomy between long GRBs connected with supernovae and short GRBs associated with kilonovae has become blurred, due to detections of long GRBs without bright supernovae and associations with kilonovae instead (Rastinejad et al., 2022; Levan et al., 2024), and short GRBs associated with supernovae (Rossi et al., 2022).
A large field-of-view infrared camera with rapid reaction to transients could quickly cover large localization areas on the sky, such as those currently provided by the Fermi/GBM, Swift/BAT, SVOM/ECLAIR and WXT onboard Einstein Probe, and detect very early counterparts when their brightness and variability amplitude is at a maximum. Observations in $K_{dark}$ are particularly relevant for very high-redshift events at $z>6$, and for those GRBs that occur deep within dusty regions (see Fig. 18). In both cases, the optical light is suppressed, and IR observations are needed to achieve detections. These observations will address critical aspects of GRB science within early observations, such as the interplay of initial emission components (internal versus external shocks; formation of forward and reverse shock; impact and interaction with circumburst medium). At late epochs, the late, non-relativistic transition can be followed, and the unobscured star-formation rate of the host galaxy estimated. Overall, the IR observations provided by Cryoscope will allow us to construct GRB afterglow samples with minimal biases due to extinction and redshift, factors that have strongly affected previous GRB observations.
Figure 18: Estimated magnitudes for GRB afterglows in $K_{dark}$ magnitudes derived from the sample of Kann et al. (2011) assuming a spectral slope of $F_{\nu}$ $\sim 1/\nu$. These light curves are corrected for extinction, which gives us a more realistic view of what is to be expected in $K_{dark}$, where the effect of extinction is less significant. We plot the 5-sigma detection limits for 15 minute exposures with the Pathfinder and the full Cryoscope. Our goal is to aim for a start of the GRB observations within 30 minutes to an hour of the alert with Cryoscope Pathfinder, and within 5 minutes of the alert for the full Cryoscope.
@@ 4.5 Solar and Interstellar Comets and Asteroids
Small bodies (asteroids and comets) are the remnants of the blocks that accreted to form the planets 4.6  Gy ago. Their importance as witnesses of solar system history has emerged in past decades, but their current description is insufficient to guide theoretical work: current observations cannot discriminate between mutually excluding hypotheses (such as a densely populated or an empty primordial asteroid belt, for instance: Morbidelli et al., 2009; Raymond & Izidoro, 2017). Models may accurately reproduce the dynamical structure of the asteroid belt, but they diverge in their prediction of the distribution of composition. It is therefore crucial to identify which compositions were native to the asteroid belt between Mars and Jupiter and which were implanted later, and what were the implantation mechanisms.
Compositional characterization of a large sample of asteroids can be achieved through broad-band spectro-photometry, complemented with albedo (DeMeo & Carry, 2013, 2014). In these respects, the upcoming NASA NEO Surveyor mission (Mainzer et al., 2023) and the Legacy Survey of Space and Time (LSST) of the Vera C. Rubin Observatory (LSST Collaboration et al., 2009; Jones et al., 2009) will play major roles—the former to measure albedo, and the latter to build multiple colors with its six visible-wavelength filters. NEO Surveyor will build upon the discoveries of WISE/NEOWISE (Wright et al., 2010; Mainzer et al., 2011), which detected $\sim 3000$ NEOs using its 3.4 and 4.6 $\,\mu$m imaging channels, enabling radiometric modeling of physical properties such as diameter and visible geometric albedo. For optical spectro-photometry with LSST, however, compositional classes of asteroids present a similar spectrum in the visible range and only differ in the infrared (DeMeo et al., 2009; Carry, 2018; Mahlke et al., 2022). By acquiring photometry in $K$-band at multiple epochs/geometries over several years, Cryoscope Pathfinder will provide for the first time the absolute magnitude and phase function (Muinonen et al., 2010) of asteroids in the near-infrared, in a similar fashion as LSST in the visible (the only way to determine the colors of small bodies with LSST, see Mahlke et al., 2021).
The difference of absolute magnitude in the visible and near-infrared will break the spectral degeneracy inherent to only visible observations, providing the compositional distribution of asteroids. Obtaining phase functions in the near-infrared also opens up new prospects for compositional mapping (Mahlke et al., 2022): these functions are known to be wavelength-dependent (Sanchez et al., 2012) but are seldom acquired with a filter other than Johnson V.
@@
In addition to the characterization of a sample of asteroids, Cryoscope Pathfinder will be able to discover new asteroids and comets which happen to be in its field-of-view during any scientific observation. Owing to its location in Antarctica, Cryoscope’s footprint will be located at high ecliptic latitude, where a deficit of detection of small bodies has been reported (Mahlke et al., 2018; Carry et al., 2021). Newly discovered asteroids of high interest, especially Near-Earth Objects, may be followed up by the Cryoscope or other facilities for improved orbit determination. Among them, objects deemed of particular importance (e.g., potential impactors) could be observed by JWST for precise measurements of size, orbit, and physical characteristics (Burdanov et al., 2025; Müller et al., 2023; Rivkin et al., 2025). However generally, if an asteroid was observed during only a short, one-night arc and was not followed up afterwards, its orbit determination will be poor. Such objects might be lost completely, or until future recovery observations. These one-night arc observations will mostly populate the Isolated Tracklet File of the Minor Planet Center. Such observations might be linked with previous or future observations done by other facilities and will contribute to a pool of objects with well-established orbits. Objects which happen to be in the field-of-view for several consecutive nights might offer better prospects of orbit determination (typically, observations from three different nights are needed for robust orbit determination of a main-belt asteroid, see Holman et al., 2018; Le Montagner et al., 2023, and references therein). Such objects will also be better placed for detection by our synthetic tracking pipeline as more images can be stacked together, increasing the objects’ S/N and enabling probes of more distant populations of small solar system bodies fainter than 20  mag (Burdanov et al., 2023).
@@ 4.6 Stellar Microlensing
Cryoscope will occupy a unique niche for gravitational microlensing experiments to detect and characterize populations of dark compact objects within our Galaxy, including primordial black holes and wide-orbit or free-floating planets. As a massive, compact object passes between us and a background star, the effect of gravitational lensing displaces, magnifies, and distorts the observed image (Paczynski, 1986). In microlensing, the displacement of the images is on the micro-arcsecond scale, and the image is effectively unresolved. However, the magnification is significant and detectable from a time-series light curve of the lensed source. The amplitude and duration of magnification depend on the geometry and relative motions of the lens-source-observer system. Still, the microlensing timescale is directly proportional to lens mass for a fixed geometry.
With the opportunity to monitor very large stellar fields, Cryoscope is particularly well-suited to microlensing experiments to conduct a gravitational census of compact objects in the Galactic disk and halo. In particular, with fast-cadence monitoring geared for minute- to day-scale microlensing events, Cryoscope is uniquely suited to probe the mass range $\sim 10^{-12}-10^{-3}$ M, from asteroid- to Jupiter-scale masses. We detail two science cases which will benefit from Cryoscope’s survey:
Wide Orbit and Free-Floating Exoplanets: Microlensing is a unique probe of ‘wide orbit’ exoplanets and especially free-floating planets (FFPs), which are exoplanets that have been ejected from their birth systems through either planet–star or planet–planet interactions (Mróz et al., 2019). FFP detections are essential for understanding the evolution of planetary systems and determining the ratio of ejected planets to bound ones (Bachelet et al., 2022) as a constraint on star- and planet-formation models. KMTNet (Jung et al., 2024; Yang et al., 2023), and MOA (Sum, 2011) have revealed a substantial population of icy exoplanets and FFPs, with hundreds of detections recorded. Modeling from the 9 years of MOA II data suggests that $19^{+23}_{-13}$ FFPs exist per star (Sumi et al., 2023), with masses ranging from terrestrial to Jovian. With its vast field of view and fast integration times, Cryoscope will enable fast-cadence exoplanet microlensing experiments to constrain the number and mass distribution of FFPs in the Milky Way.
Primordial Black Holes: In the same way as for exoplanets, Cyroscope has unique potential for microlensing searches for Primordial Black Holes (PBHs). There is a long-standing hypothesis that a cosmologically significant PBH population may be formed from critically collapsing density perturbations in the pre-inflationary universe (Zel’dovich & Novikov, 1966; Hawking, 1971). Over the last ten years, spurred partly by the possibility that PBHs may explain the unexpectedly large progenitor masses of the LIGO-Virgo mergers (Bird et al., 2016), interest in PBH formation scenarios and possible detection experiments has grown rapidly. Such a population would also be a plausible and natural solution to the dark matter problem, not requiring new particle physics beyond the Standard Model. While there exist some observational constraints on the monochromatic (i.e., uniform mass) PBH density (e.g., Carr et al., 2024, and references therein) from microlensing null results, Cryoscope represents an opportunity to determine and/or constrain the Galactic PBH population spanning the asteroid- to Jupiter-mass range, through fast-cadence ‘sit and stare’ monitoring of dark-matter-dominated sight lines towards M31 (e.g. Niikura et al., 2019; Blaineau et al., 2022) or the LMC (Mróz et al., 2024).
Cryoscope’s strengths: For the microlensing science cases above, the science-limiting factors are: (i) a vast field of view to allow simultaneous monitoring of the large number of stars; (ii) a good and stable atmospheric seeing and transparency to maximize sensitivity to microlensing amplifications and minimize blending and/or contamination; (iii) fast integrations and read-out for efficient high-cadence monitoring. Cryoscope delivers on all counts, with the additional benefit of operating in the NIR, where Galactic dust obscuration/attenuation is minimal (e.g., Shvartzvald et al., 2017, 2018; Kondo et al., 2023), and the benefit of very-long night observing for maximally efficient searches on few-hour timescales.
Relative to mega-projects like Rubin, Euclid, and Roman, Cryoscope has the opportunity to conduct repeat ‘sit-and-stare’ microlensing experiments over multiple Galactic and extragalactic sight lines. Cryoscope’s ability to trace such regions of high stellar density will increase the probability of detecting multiple microlensing signals, and provide better event statistics to distinguish between the kinematics of baryonic (planetary) and non-baryonic (PBH) microlenses. With the combination of space-competitive seeing and a 4.3$\times$ advantage over DECam in field-of-view, Cryoscope can establish and exploit a unique technical and scientific niche in microlensing experiments, even in the era of Rubin, Euclid, and Roman.
@@ 4.7 Core Collapse Supernovae
Massive stars with $\gtrsim 8~{}M_{\odot}$ form an iron core towards the end of their life and die when this iron core exceeds the Chandrasekhar mass and collapses, sending out a shock wave that disrupts the star. The resulting core-collapse supernova (CCSN) forms a diverse group of transients whose properties are determined by the properties of the progenitor star. For instance, the amount of hydrogen left in the envelope determines the spectral type of the CCSN (Filippenko, 1997); the star’s mass at the time of explosion determines how long the light curve takes to rise (Khatami & Kasen, 2019); and the core mass of the star determines the strengths of different emission lines later in the nebular phase (e.g., Jerkstrand et al., 2014). While most of what we know about CCSNe as a class have been derived from optical observations, increasing numbers of IR observations in the past decade revealed several aspects of massive stellar explosions that had remained elusive. Cryoscope observations will help us probe the chemical evolution of CCSNe ejecta, reveal late-time power sources of CCSNe, and uncover a population of CCSNe hidden behind dust.
@@
While JWST observations are extremely powerful in tracking molecule and dust formation in CCSNe, the number of observations is still limited to a few nearby and notable objects (e.g., Hosseinzadeh et al., 2023; Shahbandeh et al., 2023). Cryoscope is unique as the only planned ground-based, high-cadence survey that covers wavelengths beyond 2 $\,\mu$m (unlike the DREAMS and WINTER surveys powered by InGaAs detectors). The $K_{dark}$ band, unlike the more typical $K_{s}$, covers the bandhead of CO’s first overtone band at around 2.3 $\,\mu$m. $K_{dark}$ observations will also be able to track the rising thermal continuum from newly formed dust grains in the ejecta of CCSNe. Unlike in previous targeted IR observations, Cryoscope will observe a statistically significant sample of CCSNe and measure the evolution of CO and dust formation to test chemical evolution models of CCSNe.